ALLOWANCE FOR DOUBTFUL DEBTS Zet Ltd. decided to make allowance for doubtful debts at the rate of 5% of trade receivables at the end of its financial year, that is 31 December each year. He made the first allowance on 31 December 2017. Trade receivables has been as follows: Year 2017 2018 2019 2020 Trade Receivables (£) 114,00 12,200 10,800 11,200. prepare journal entries Suggestion It should look like this Bad debt Expense xxxxxx Debit Provision for Doubtful DebtsxxxxCredit
Answer #1

Year 2017:

Bad debt expense = Unadjusted trade receivable × 5%

                              = 11,400 × 5%

                              = 570

Journal

Date

Account titles and explanation

Ref.

Debit

Credit

31/12/17

Bad debt expense

570

          Provision for doubtful debts

570

To record of making allowance for doubtful debts.

Year 2018:

Target balance = Unadjusted trade receivable × 5%

                        = 12,200 × 5%

                        = 610

Bad debt expense = Target balance – Unadjusted credit balance on provision for doubtful debts

                              = 610 – 570

                              = 40

Journal

Date

Account titles and explanation

Ref.

Debit

Credit

31/12/18

Bad debt expense

40

          Provision for doubtful debts

40

To record of making allowance for doubtful debts.

Year 2019:

Target balance = Unadjusted trade receivable × 5%

                        = 10,800 × 5%

                        = 540

Unadjusted credit balance on provision for doubtful debts = 570 + 40 = 610

Since target balance is lower than the unadjusted credit balance on provision for doubtful debts, there should be a written of entry.

Written-off amount = Unadjusted credit balance on provision for doubtful debts – Target balance

                                = 610 – 540

                                = 70

Journal

Date

Account titles and explanation

Ref.

Debit

Credit

31/12/19

Provision for doubtful debts

70

        Accounts receivable

70

To record written off an uncollectible account.

Note: there should not be any bad debt account in this year.

Year 2020:

Unadjusted trade receivable = 11,200 – 70 = 11,130

Target balance = Unadjusted trade receivable × 5%

                        = 11,130 × 5%

                        = 556.50

Unadjusted credit balance on provision for doubtful debts = 610 – 70 = 540

Bad debt expense = Target balance – Unadjusted credit balance on provision for doubtful debts

                              = 556.50 – 540

                              = 16.5

                              = 17 rounded

Journal

Date

Account titles and explanation

Ref.

Debit

Credit

31/12/20

Bad debt expense

17

          Provision for doubtful debts

17

To record of making allowance for doubtful debts.